Grade 11Mechanics

an object of mass 1 kgmoves under the action of a force F in a straight line with its velocity V , changing with displacement x as v =2rootx. find the work done in a displacement from x=0 to x=2.
(in joule)

vikki kumar
8 Years agoGrade 11

1 Answer

Bitanuka
8 Years ago
M = 1kg, Force = FTherefore, W= F*x = F(2-0) = 2FF = dp/dt = d(mv)/dt = m*dv/dt = d(2√x)/dt = 2*d(√x)/dt =2*1/2*1/(√x) = 1/(√x) = 1/√2Therefore, W = 2*1/√2 = √2